How many months will it atke to pay off a credit card with a balance of of $1000 if you only pay the minimum payment of $20 each month? Your Annual Percentage Rate is 12.96%
Answer by ewatrrr(24785)


Hi, previousy posted
Note: formula for N(the number of months)to pay off the $1000 is as follows:
N = −log(1−i*A/P) / log(1+i) where A = 1000, P = 20 and i = 12.96/12 = 1.08/mo
N = −log(1−(.0108)*1000/20) / log(1.0108)
N = −log(1−.54)/ log(1.0108)
N = - log (.46) / log(1.0108) = -(-.7765)/.0107 = 72.57 months
After 72 months there will be a balance which can be determined by using:
B = A(1+i)^72 - P/i[(1+i)^72 - 1]
B = 1000(1.0108)^72 - 20/.0108[(1.0108)^72 - 1]